OpenStudy (desmarie):

Simplify the following expression. rad3 times 2rad2

OpenStudy (jclark):

Have you ever seen this notation?\[x^{\frac{ 1 }{ 3 }}\times x ^{\frac{ 1 }{ 3 }}\]

OpenStudy (desmarie):

the options they give to choose an answer from are: 6rad2 rad12 6rad5 2rad6

OpenStudy (desmarie):

no I dont thinks so.....

OpenStudy (desmarie):

*think

OpenStudy (jclark):

here's the general rule: \[\sqrt{x }\times \sqrt{y} =\sqrt{xy}\]

OpenStudy (jclark):

so the whole thing reads \[\sqrt{3} \times 2 \times \sqrt{2}\]

OpenStudy (desmarie):

ok..i am following you so far :)

OpenStudy (jclark):

you can multiply things in any order, so it could be \[2 \times \sqrt{3} \times \sqrt{2}\]

OpenStudy (jclark):

now use the general rule I showed you. What should the radical be now?

OpenStudy (jclark):

hint: x is 3 and y is 2

OpenStudy (desmarie):

sorry.....the answer is 2rad6....right?

OpenStudy (jclark):

if x=3 and y=2 then \[\sqrt{x} \times \sqrt {y} = \sqrt {3} \times \sqrt{2} = \sqrt {xy} = \sqrt {3 \times 2}\]

OpenStudy (jclark):

you tell me...

OpenStudy (desmarie):

I would say 2rad6

OpenStudy (jclark):

And you would be correct.

OpenStudy (desmarie):

thanks so much for your help